Assume Binomial Distribution with n = 10 and p = 0.6. Please show your answers to 4 decimal places.

P(X = 4) =  

P(X ≤ 4) =   

P(X < 4) = 

P(X >4) =   

P(X ≥ 4) = 


1
Expert's answer
2022-03-19T08:53:46-0400

"n=10 ~~p=0.6~~q=1-p=0.4"


(a) "P(X=4)"

"=\binom{n}{x}\times p^xq^{n-x}"


"=\binom{10}{4}\times0.6^40.4^6"


"=0.1115"


(b) "P(X\le4)"

"=P(4)+P(3)+P(2)+P(1)+P(0)"

"=0.1115+0.0425+0.0106+0.0016+0.0001"


"=0.1663"


(c) "P(X<4)"

"=P(3)+P(2)+P(1)+P(0)"

"=0.0425+0.0106+0.0016+0.0001"

"=0.0548"


(c) "P(X>4)"

"=1-P(X\le4)"

"=1-0.1663"

"=0.8337"